OpenStudy (anonymous):

Auditions for a new, upcoming band in Chicago were held with 124 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 28, how many women auditioned?

OpenStudy (alexwee123):

ugh linear algebra o.0

OpenStudy (alexwee123):

so you can get 2 equations from the info but first lets make x=men and y=women because 124 people in total signed up we can get an equation of x+y=124 and from " If twice the number of men less three times the number of women is equal to 28" we can get 2x-3y=28

OpenStudy (alexwee123):

so use substitution or elimination to find the # of women

OpenStudy (anonymous):

so what do i substitute into 2x-3y=28?

OpenStudy (alexwee123):

either x=124-y or y=124-x but i would use x=124-y because you're trying to find the number of women

OpenStudy (anonymous):

248-5y=28??

OpenStudy (alexwee123):

yep :D

OpenStudy (anonymous):

5y=256?

OpenStudy (anonymous):

y=128 thats the answer?

OpenStudy (anonymous):

no that cant be the answer.. the choices are. 40,44,70, or 80.

OpenStudy (alexwee123):

248-5y=28? -5y=-220 y=44

OpenStudy (anonymous):

ohhh true!